# Incorporating Procedural Fairness in Flag Submissions on Social Media Platforms

### v1

@@ 2.3. Procedural Fairness in Content Moderation
Fairness is a multifaceted concept that includes procedural fairness (Vaccaro et al., 2020; Lee et al., 2019), outcome fairness (Ma et al., 2023a; Vaccaro et al., 2020), and restorative fairness (Kou, 2021; Schoenebeck et al., 2021), emphasizing process, outcome, and feedback, respectively. Scholars focused on incorporating fairness in decision-making systems highlight enacting unbiased processes or outcomes, avoiding unjust situations (Shin and Park, 2019), and treating every individual with respect, dignity (Katsaros et al., 2024), and equality (Ma and Kou, 2022). Content moderation scholars have argued that achieving fairness in moderation processes entails administering equitable treatment of all users (Jhaver et al., 2019a) through clear and consistent criteria for content review that integrate the needs of marginalized groups, such as racial minorities and LGBTQ+ individuals (Haimson et al., 2021; Ma and Kou, 2022).
@@ 3. Components of Flagging Design
Our study examines four components of flagging mechanisms commonly adopted by social media platforms, each characterized by varying levels of user engagement. Drawing on the characteristics of each component and insights from previous literature, we formulate hypotheses about how these factors contribute to different fairness attributes.
3.1. Flag Classification Levels
As users engage in the flagging process, they may encounter a classification scheme with primary menus and submenus (see Figure 2) that specify the range of rule violations the flagged post may have, such as hate speech and spam. Making selections from predefined rule violations lets users categorize their concerns with the flagged post (Crawford and Gillespie, 2016). From a platform perspective, imposing this precisely designed classification scheme on users is crucial to triage flagged posts, validate whether the selected rule violation occurred, and highlight that the content of flagged posts drives flag review decisions (instead of other factors, such as flaggers’ identity). However, it is not yet clear how the design of classification schemes affects flaggers’ perceptions of fairness.
@@ 3.1. Flag Classification Levels
We expect that requiring users to specify the violation within flagged content with higher granularity will prompt them to pay closer attention to the categories themselves and classify the rule violation categories cautiously. Since the rule violation scheme is derived from content moderation guidelines established by the platform, users may perceive that the moderation process relies on clear and precise rules when the classification scheme accommodates more specific rule violation selections. Further, incorporating more comprehensive categories into the flagging mechanism may influence users to perceive that the moderation criteria are consistent rather than arbitrary or situational. Therefore, we expect that if the flagging mechanism offers users more detailed categories for flagging, they will perceive that the system prioritizes its formalized moderation criteria rather than other factors, such as flagger identity.
@@
Integrating the rule violation classification into the flagging mechanism serves a dual purpose: (a) it clarifies the reporting rationale from the user’s perspective and (b) it reminds users of the platform’s declared categories of rule violations. Offering users a more detailed classification of rule violations deepens their comprehension of the types of content subject to moderation. Since transparency is linked to the amount of information about the process  (Lee et al., 2019; Rader et al., 2018), furnishing a more detailed rule violation classification scheme to users may enhance the transparency of the flagging review process. Therefore, we hypothesize:
@@
Providing a more detailed rule violation classification scheme means that users can express their opinions more precisely on the platform. In other words, allowing users to select both a primary category and a subcategory sequentially ensures a more precise conveyance of their intentions during the review process compared to having no or only a single-level classification scheme option. Given that voice is closely tied to providing opportunities for users’ perspectives to be expressed within the process, incorporating more detailed levels of rule violation classification scheme may lead users to perceive that the platform has a genuine interest in understanding their concerns, which may result in an elevated perception of having a voice in the process. We therefore hypothesize:
@@
3.2. Posting Guidelines
Today, all major social media platforms publicly list their guidelines for appropriate conduct or terms of use on their sites. These guidelines are crucial for designing accountable content moderation systems because they serve as the official criteria for the platform’s moderation decisions (Gillespie, 2018a; Zolides, 2021; Maddox and Malson, 2020). Prior research also shows that drawing users’ attention to posting guidelines contributes to pro-social community outcomes. For example, Matias (2019b) demonstrated through an online experiment on Reddit that announcing posting guidelines in public discussions increases the likelihood of compliance with them and encourages newcomer participation. Bhandari et al. (2021) show that users who look at the posting guidelines are more likely to flag harassment posts. Further, when guidelines are easily noticeable, users are more likely to accept the platform’s moderation decisions  (Bradford et al., 2019; Kiesler et al., 2012) and perceive them as fair  (Jhaver et al., 2019a).
While this literature has established the benefits of how users’ attention to posting guidelines enhances their fairness perception of moderation outcomes, little attention has been given to integrating these guidelines into flagging mechanisms and exploring its impact. Further, Zhang et al. (2023) noted deficiencies in the clarity and details of how rule violations are defined in the current social media reporting systems. Therefore, we posit that enhancing guideline visibility and details may improve user perceptions about the flagging process. We investigate how different granularity levels of posting guidelines influence users’ perceptions, specifically regarding the fairness of the flagging process.
Guidelines function as the rubric for content moderation, lending greater legitimacy to the review process (Crawford and Gillespie, 2016). Offering clear posting guidelines in the flagging mechanism may lead users to perceive that the review process adheres to a predetermined rubric, suggesting that the flagged content’s review process is largely influenced by the specified guidelines and not other external factors. Therefore, the visibility of guidelines in this process can augment users’ perception of consistent review procedures grounded in established principles. Additionally, we maintain that more detailed guidelines, presented with specific examples of rule violations, could strengthen the perception of a robust rule-based review process. Therefore, we hypothesize:
Integrating posting guidelines in the flagging mechanism is a way of disclosing content review criteria by illustrating possible flaggable posts, thereby enhancing users’ understanding of the content review process (Ozanne et al., 2022). Thus, more detailed guidelines, including rule-violation example posts, will give users more information on the review criteria, thereby promoting greater transparency. We therefore hypothesize that:
Providing guidelines for posting in the flagging procedure may impose a sense of the need for strict adherence to predefined rules. Since users sometimes use flagging in ways that may not align with platforms’ notion of flaggability (Kou and Gui, 2021), e.g., to express their objections to others’ opinions (Gillespie, 2018a), requiring them to follow the guidelines may discourage users from flagging objections that have not yet been formalized as rule violations in the guidelines. Therefore, when a flagging mechanism prompts users to assess flaggability by providing clear posting guidelines, they may perceive a lack of opportunities to express their opinions freely. We thus hypothesize:
3.3. Text Box
From the perspective of end-users, most content moderation processes have rigid protocols that offer limited opportunities for expression or interaction with platform administrators (Jhaver et al., 2023; Shahid and Vashistha, 2023). However, some mechanisms offer flexibility, such as enabling users to express their opinions freely in their own words. One example is the text box in a flagging mechanism that could invite users to express their opinions before a flagging decision is made. Flaggers could clarify their rationale for reporting by providing relevant context and detailing the rule violations that they believe have occurred. Although Wilson and Land (2020) found that the flagging mechanism can enhance users’ trust in moderation decisions by offering opportunities to explain the reasons for reporting, such as the post’s context, most social media platforms do not let users express their reasons for flagging a post in an open-ended manner.
We examine the utility of offering such a text box in the flagging mechanism, enabling users to explain their reasons for flagging, including how toxic posts impact them. We expect that this text box would empower flaggers to actively participate in content moderation by detailing their objections, and its unrestricted nature would let them express their opinions in diverse and personalized ways.
By detailing the context of any flagged post and why it is inappropriate in a text box, flaggers provide moderators with a basis for post review. The availability of this text box may signal to users that moderators incorporate this input into their decisions — this offers additional insight into the review process. It may reinforce users’ perceptions that their input strengthens the justification for the flag review and its outcome. Thus, we expect that providing users with a text box would enhance their perception of transparency about the review of flagged content, and we thus hypothesize:
Including a text box in the flagging process lets users articulate their reasons for flagging with higher precision. This feature empowers users by granting them substantial autonomy to express their flagging intentions beyond the confines of predefined options provided by the platform’s classification scheme. By enabling users to elaborate on their opinions rather than relying solely on predetermined categories of inappropriate content, this approach addresses the limitations of the classification scheme. Consequently, providing a text box may lead users to perceive that the flagging mechanism values their opinions and intentions more than when they cannot express their views in their own words. Based on this, we expect that users will have a higher perception of voice when the flagging mechanism offers a text box, and we hypothesize:
3.4. Moderator Type
After users submit a flag for moderation, it undergoes assessment by either a human or automated moderator (Pan et al., 2022; Seering, 2020) who evaluates whether the flagged content should be kept on the site or sanctioned. Platforms hire either third-party contractors to serve as commercial content moderators (Gillespie, 2018a; Roberts, 2016) or rely on end-users to serve as volunteer moderators (Jhaver et al., 2019b; Matias, 2019a) to review flagged posts. Automated mechanisms to evaluate flagged posts typically use hashing technology or natural language processing tools to identify rule violations in the posted text, images, and videos (Chandrasekharan et al., 2019; Díaz and Hecht-Felella, 2021).
While human moderators are better equipped to understand the cultural nuances and context of posts (Zhou et al., 2023), automated moderation can minimize delays by taking immediate action and addressing the challenges of scale (Chandrasekharan et al., 2019). Several studies indicate a preference among users for human decision-makers over AI counterparts due to perceptions that humans can better recognize emotions (Lyons et al., 2022; Wojcieszak et al., 2021). Further, Pan et al. (2022) found that users perceive moderation decision-making from a human expert panel as more legitimate than decisions made by paid platform reviewers or algorithms. However, some research suggests that whether a moderator is human or AI does not significantly affect perceptions of fairness (Jhaver et al., 2019a), transparency (Ozanne et al., 2022; Gonçalves et al., 2021) and trust (Molina and Sundar, 2022) in moderation decisions.
Though these prior studies explore how the choice of AI versus human decision-makers shapes user perceptions of moderation outcomes, we know relatively little about how different types of moderators impact user participatory processes like flagging. In addition, prior research has mainly focused on users’ perceptions when provided with moderation information alongside decision outcomes (Pan et al., 2022; Jhaver et al., 2019c; Molina and Sundar, 2022). In contrast, our study examines the impact of revealing moderator information before decision-making. We investigate how different scenarios in flagging mechanisms involving three types of moderator information (human, bot, not specified) affect users’ perceptions of the fairness of flag review.
Human moderators are proficient in incorporating the context and their domain knowledge when regulating posts, but they risk introducing potential biases in decision-making (Pan et al., 2022). In contrast, automated moderators make decisions based on objective, pre-defined criteria (Ozanne et al., 2022), but they may be unable to account for some relevant context. Thus, we posit that users will perceive the flag review process conducted by an automated bot as more consistent than a review conducted by a human or an unknown moderator and thus hypothesize:
Transparency in content moderation fosters a meaningful understanding of the moderation processes and improves accountability (Suzor et al., 2019). Lack of information about the moderator type may lead users to speculate about who reviews their flagged posts, thereby reducing transparency about the process (Myers West, 2018; Suzor, 2019). Therefore, we hypothesize that providing information about whether a human or a bot would review the flag can improve the perceived transparency of the flagging system compared to withholding such information.
Prior studies have shown that users perceive decisions made by human moderators, such as expert juries, to be more legitimate than those made by AI (Pan et al., 2022). Additionally, users prefer expressing their opinions to human moderators during the appeal process since they believe that humans are more likely to examine their opinions with empathy and compassion (Yurrita et al., 2023). Therefore, we hypothesize that in the context of flagging interactions, users would feel they have a greater voice when a human moderator is involved in the review process instead of an automated moderator or an unknown source.
4. Methods
We begin this section by describing how we implemented 54 distinct flagging scenarios using Qualtrics to survey our participants and explain how we operationalized our variables of interest. Next, we offer information about our survey deployment and detail our ethical considerations. Finally, we describe the methods we used to analyze the Likert scale and open-ended responses from our survey data. We preregistered this study at OSF. [^4] https://osf.io/982wd/?view_only=91e6792f777b44b582fd6bb480ab142c
4.1. Survey Design
To test how the availability and granularity of different flagging components affect users’ perceived fairness, we created an online survey using Qualtrics software. We distributed the survey via Lucid Theorem, an academically-oriented survey platform, to collect responses from U.S. adult internet users. The survey began with a page displaying seven examples of rule-violating posts, shown in Table 2. Warning: The example posts in Table 2 may be disturbing to some readers.
No. Post
     6 Omg just got tons of Bucks from here! - [xyz phishing site].com.
The flag submission choices shown to participants followed a between-subjects factorial design created by combining different implementation choices for the four flagging components. It comprised 3 (Rule violation classification scheme levels) $\times$ 3 (Granularity of posting guidelines) $\times$ 2 (Availability of a text box) $\times$ 3 (Types of moderators) different conditions. Depending on their assigned scenario, participants experienced a different design of the rule violation classification scheme, varying levels of detail regarding platform guidelines, the availability (or absence) of an open-ended text box, and information about the moderator (human, bot, or no information) of the flag review process.
We now detail how we operationalized four components of flagging designs within the Qualtrics software.
4.1.1.   Independent Variables: Components of Flagging Mechanism
(1) Rule Violation Classification Scheme
In our survey, we organized rule violation classification schemes into nine distinct primary categories, each with additional submenus, as shown in Table 3. This classification scheme was developed by referencing rule violation categories in flagging mechanisms across major social media platforms, including Facebook, Reddit, X, and YouTube. By synthesizing rule violation categories across multiple platforms, we aimed to allow users to explore flagging mechanisms without focusing on the norms for any particular platform.
Table 3. Rule Violation Classification Scheme Levels.
@@ 4.1.1.   Independent Variables: Components of Flagging Mechanism
Each category includes an option for ‘Something else.’
@@
(2) Platform Guidelines
We developed a set of platform guidelines that explicitly describe the types of content prohibited on the platform. Table 4 shows this list and their corresponding examples. We developed these guidelines by synthesizing Facebook, Reddit, X, and YouTube guidelines. We retrieved the guidelines from webpages of these sites under titles such as ‘community standards,’ ‘content policy,’ ‘rules and policy,’ or ‘community guidelines & policies.’ By identifying common guideline themes across platforms, we compiled a core list of community guidelines and incorporated specific language from different platforms to clarify them. For example, all platforms address violence-related activities, categorizing them as violence and incitement (Facebook), threats of violence and incite violence (Reddit), glorification of violence, violent and hateful entities, and violent speech (X), and violent criminal organizations and violent or graphic content (YouTube). We consolidated these various forms of violent activities into one overarching category of ‘violence,’ recognizing it as a fundamental aspect of platform guidelines.
To construct detailed guidelines, we gathered illustrative examples for each rule violation category to provide additional context. For many of these guidelines, we used examples officially offered by the platforms we considered (Facebook, Reddit, X, and YouTube) when such examples were available. For instance, YouTube provides several examples of hate speech, from which we selected and adjusted two for our guidelines. We could not find official examples for some rule violation categories in our list. In such cases, we conducted a content search on the platforms with keywords borrowed from platform guidelines to find explicit examples illustrating prohibited activities. For instance, X addresses destructive behavior with keywords such as physical injuries, cutting, eating disorders, bulimia, and anorexia, which we adopted to search for explicit examples to include in our guidelines.
Table 4. Granularity of Posting Guidelines.
Basic guidelines: ”We do not allow content that” Examples of posts that violate guidelines
@@
(3) Text Box
A text box was integrated into the flagging mechanism using the text entry feature within Qualtrics. Figure 3 displays how this box appeared during the survey. Each participant experienced one of the following two conditions:
Figure 3. Participants can express their opinions using the text entry feature provided by Qualtrics.
@@
This text box enabled participants to input their responses without any word limit, and it was furnished with the prompt, “Please describe your problems with this post. Your response will let us know what’s happening and help us review this post.
(4) Moderator Type
The last component integrated into our flagging mechanism included a thank you message to users for their flagging action, optionally accompanied by moderator information. In the first condition, participants received only a message expressing gratitude for flagging. In the second condition, the gratitude message was exhibited alongside the information that a human moderator will be responsible for reviewing the flag. Finally, in the third condition, the gratitude message noted that an automated moderator will review the flagged post. Specifically, we used the following message texts in each condition:

### v2

@@ 2.3. Procedural Fairness in Content Moderation
Fairness is a multifaceted concept that includes procedural fairness (Vaccaro et al., 2020; Lee et al., 2019), outcome fairness (Ma et al., 2023a; Vaccaro et al., 2020), and restorative fairness (Kou, 2021; Schoenebeck et al., 2021), emphasizing process, outcome, and feedback, respectively. Scholars focused on incorporating fairness in decision-making systems highlight enacting unbiased processes and outcomes, avoiding unjust situations (Shin and Park, 2019), and treating every individual with respect, dignity (Katsaros et al., 2024), and equality (Ma and Kou, 2022). Content moderation scholars have argued that achieving fairness in moderation processes entails administering equitable treatment of all users (Jhaver et al., 2019a) through clear and consistent criteria for content review that integrate the needs of marginalized groups, such as racial minorities and LGBTQ+ individuals (Haimson et al., 2021; Ma and Kou, 2022).
@@ 3. Components of Flagging Design
Our study examines four components of flagging mechanisms commonly adopted by social media platforms. Drawing on the characteristics of each component and insights from previous literature, we formulate hypotheses about how these factors contribute to different fairness attributes in this section. We also describe how we operationalized each component through our survey questions.
@@
3.1. Flag Classification Levels
As users engage in the flagging process, they may encounter a classification scheme with primary menus and submenus (see Figure 2) that specify the range of rule violations the flagged post may have, such as hate speech and spam. Making selections from predefined rule violations lets users categorize their concerns with the flagged post (Crawford and Gillespie, 2016). From platforms’ perspective, imposing this precisely designed classification scheme on users is crucial to triage flagged posts, validate whether the selected rule violation occurred, and highlight that the content of flagged posts drives flag review decisions (instead of other factors, such as flaggers’ identity). However, it is not yet clear how the design of classification schemes affects flaggers’ perceptions of fairness.
@@ 3.1. Flag Classification Levels
Since the rule violation scheme is derived from content moderation guidelines, we expect that if the flagging mechanism offers users more detailed categories for flagging, they will perceive that the system prioritizes its formalized moderation criteria rather than other factors, such as flagger identity. Therefore, we hypothesize:
@@
Transparency is linked to the amount of information about the process  (Lee et al., 2019; Rader et al., 2018). Thus, furnishing a more detailed rule violation classification scheme may enhance the transparency of the flag review process. Therefore, we hypothesize:
@@
Given that voice is closely tied to providing opportunities for users’ perspectives to be expressed within the process (Katsaros et al., 2024, 2022a), incorporating more detailed levels of rule violation classification scheme may lead users to perceive that the platform has a genuine interest in understanding their concerns, which may result in an elevated perception of having a voice in the process. We therefore hypothesize:
@@
Table 2. A taxonomy of rule violation categories developed by referencing flag interfaces across multiple platforms. We used this taxonomy to operationalize flag classification choices in our survey. Each category below also included an additional submenu option of ‘Something else.’
@@
Operationalization
In our survey, we organized rule violations into nine distinct primary categories, each with an additional submenu, as shown in Table 2. This classification scheme was developed by referencing rule violation categories in flagging mechanisms across major social media platforms, including Facebook, Reddit, X, and YouTube. By synthesizing rule violation categories across multiple platforms, we aimed to allow users to explore flagging mechanisms without focusing on the norms for any particular platform.
@@
3.2. Posting Guidelines
Posting guidelines are crucial for designing accountable content moderation systems because they serve as the official criteria for the platform’s moderation decisions (Gillespie, 2018a; Zolides, 2021; Maddox and Malson, 2020). Prior research also shows that drawing users’ attention to posting guidelines contributes to pro-social community outcomes. For example, Matias (2019) demonstrated through an online experiment on Reddit that announcing posting guidelines in public discussions increases the likelihood of compliance with them and encourages newcomer participation. Further, when guidelines are easily noticeable, users are more likely to accept the platform’s moderation decisions  (Bradford et al., 2019; Kiesler et al., 2012) and perceive them as fair  (Jhaver et al., 2019a).
While this literature has established the benefits of how users’ attention to posting guidelines enhances their fairness perception of moderation outcomes, little attention has been given to integrating these guidelines into flagging mechanisms and exploring its impact. We investigate how different granularity levels of posting guidelines influence users’ perceptions, specifically regarding the fairness of the flagging process.
Offering clear posting guidelines in the flagging mechanism may lead users to perceive that the review process adheres to a predetermined rubric and is largely influenced by the specified guidelines and not other external factors. Additionally, we posit that more detailed guidelines, presented with specific examples of rule violations, could strengthen the perception of a robust rule-based review process. Therefore, we hypothesize:
Integrating posting guidelines in the flagging mechanism is a way of disclosing content review criteria by illustrating possible flaggable posts, thereby enhancing users’ understanding of the content review process (Ozanne et al., 2022). We therefore hypothesize that:
Showing guidelines in the flagging procedure may impose a sense of the need for strict adherence to predefined rules. Since users sometimes flag in ways that may not align with platforms’ notion of flaggability (Kou and Gui, 2021), requiring them to follow the guidelines may lead to perceptions that they cannot express their opinions freely. We thus hypothesize:
Operationalization
We developed a set of platform guidelines that explicitly describe the types of content prohibited on the platform. Table 3 shows this list and their corresponding examples. We developed these guidelines by synthesizing Facebook, Reddit, X, and YouTube guidelines. We retrieved the guidelines from webpages of these sites under titles such as ‘community standards,’ ‘content policy,’ ‘rules and policy,’ or ‘community guidelines & policies.’ By identifying common guideline themes across platforms, we compiled a core list of community guidelines and incorporated specific language from different platforms to clarify them.
To construct detailed guidelines, we gathered illustrative examples for each rule violation category through two means: (1) using examples officially offered in platforms’ description of community guidelines and (2) conducting a content search on platforms with keywords borrowed from guidelines.
Table 3. Posting guidelines developed by synthesizing guidelines across Facebook, Reddit, X, and YouTube and used in our survey interface. A prompt of ‘We do not allow content that’ appeared below these guidelines were shown.
     Posting Guidelines Examples of Posts Violating the Guideline
@@ 3.2. Posting Guidelines
3.3. Text Box
From the perspective of end-users, most content moderation processes have rigid protocols that offer limited opportunities for expression or interaction with platform administrators (Jhaver et al., 2023; Shahid and Vashistha, 2023). However, some mechanisms offer flexibility e.g., a text box in a flagging mechanism that invites users to express their opinions before a flagging decision is made. We examine the utility of offering such a text box in the flagging mechanism, enabling users to explain their reasons for flagging, including how toxic posts impact them. We expect that this text box would empower flaggers to actively participate in content moderation by detailing their objections, and its unrestricted nature would let them express their opinions in diverse and personalized ways.
By detailing the context of any flagged post and why it is inappropriate in a text box, flaggers provide moderators with a basis for post review. The availability of this text box may signal to users that moderators incorporate this input into their decisions — this offers additional insight into the review process. We thus hypothesize:
Including a text box in the flagging process lets users articulate their reasons for flagging with higher precision. This feature empowers users by granting them substantial autonomy to express their flagging intentions beyond the confines of predefined options provided by the platform’s classification scheme. We thus hypothesize:
Operationalization
We integrated a text box in our survey’s flagging mechanism using the text entry feature provided by Qualtrics. This text box enabled participants to input their responses without any word limit, and it was furnished with the prompt, “Please describe your problems with this post. Your response will let us know what’s happening and help us review this post.” Figure 3 displays how this box appeared during the survey. Each participant experienced one of the following two conditions:
Figure 3. Participants can express their opinions about the flagged post using the text entry feature provided by Qualtrics.
@@ 3.3. Text Box
3.4. Moderator Type
After users submit a flag for moderation, it undergoes assessment by either a human or automated moderator (Pan et al., 2022; Seering, 2020) who evaluates whether the flagged content should be kept on the site or sanctioned. Several studies indicate a preference among users for human decision-makers over AI counterparts due to perceptions that humans can better recognize emotions (Lyons et al., 2022; Wojcieszak et al., 2021). However, some research suggests that whether a moderator is human or AI does not significantly affect perceptions of fairness (Jhaver et al., 2019a), transparency (Ozanne et al., 2022; Gonçalves et al., 2021) and trust (Molina and Sundar, 2022) in moderation decisions.
Though these prior studies explore how the choice of AI versus human decision-makers shapes user perceptions of moderation outcomes, we know relatively little about how different types of moderators impact user participatory processes like flagging. We investigate how different scenarios in flagging mechanisms involving three types of moderator information (human, bot, not specified) affect users’ perceptions of the fairness of flag review.
Human moderators are proficient in incorporating the context and their domain knowledge when regulating posts, but they risk introducing potential biases in decision-making (Pan et al., 2022). In contrast, automated moderators make decisions based on objective, pre-defined criteria (Ozanne et al., 2022), even though they may be unable to account for some relevant context. Thus, we hypothesize:
Transparency in content moderation fosters a meaningful understanding of the moderation processes and improves accountability (Suzor et al., 2019). Lack of information about the moderator type may lead users to speculate about who reviews their flagged posts, thereby reducing transparency about the process (Myers West, 2018; Suzor, 2019). Therefore, we hypothesize:
Prior studies have shown that users perceive decisions made by human moderators, such as expert juries, to be more legitimate than those made by AI (Pan et al., 2022). Additionally, users prefer expressing their opinions to human moderators during the appeal process since they believe that humans are more likely to examine their opinions with empathy and compassion (Yurrita et al., 2023). We thus hypothesize:
Operationalization
Following the other flagging prompts, we included a thank you message to users for their flagging action in our survey, optionally accompanied by moderator information. Specifically, each participant saw one of the following messages: